OpenStudy (anonymous):

place the answer in scentific notation, when appropriate and simplify the units. (1.33)sin35.0=(1.50)sinθ 35.0 is in degrees

OpenStudy (anonymous):

\[\left( 1.33 \right)\sin 35.0=\left( 1.50 \right)\sin \theta \]

OpenStudy (anonymous):

35.0 is in degrees